# Robust Preference Optimization: Aligning Language Models with Noisy Preference Feedback

### v1

@@ 5.1 Experimental setup
All experiments were conducted using the Hugging Face TRL library (von Werra et al., 2020). We performed full-parameter fine-tuning for all models, leveraging DeepSpeed ZeRO-3 (Rasley et al., 2020) to manage memory across our hardware setup. Models were trained for a single epoch using the AdamW optimizer (Loshchilov and Hutter, 2019) with a cosine learning rate scheduler, a peak learning rate of $5\times 10^{-7}$, and a global batch size of 64. We also maintained an Exponential Moving Average (EMA) of the model parameters with a decay rate of $\alpha=0.1$. The temperature for DPO and its variants was set to $\beta=0.01$. For all RPO variants, the reliability parameter was initialized to $\eta^{(0)}=0.9$ and was updated via the E-step after the training epoch was complete. All training was performed on a cluster of 8x NVIDIA A800 40GB GPUs.
@@
To demonstrate that RPO operates as a versatile meta-framework, we benchmark it against four popular direct preference alignment methods: DPO (Rafailov et al., 2023); IPO (Azar et al., 2023), which uses a squared hinge loss to optimize preferences; SimPO (Meng et al., 2024), which proposes a simplified, reference-free reward formulation normalized by sequence length; and CPO (Xu et al., 2024), which adds a term to directly maximize the likelihood of the preferred response. For each of these baselines, whose loss functions are detailed in Table 1, we compare the original algorithm to its RPO-enhanced counterpart (e.g., DPO vs. R-DPO).
@@ 5.2 Main results
As shown in Table 2, our experimental results provide strong evidence for the effectiveness of the RPO framework. We summarize the key findings below.
RPO as a meta-framework.
The primary takeaway from our experiments is that RPO functions as a highly effective and generalizable meta-framework. Across all four baseline algorithms (DPO, IPO, SimPO, CPO), both base models (Mistral-7B and Llama-3-8B), and all evaluation benchmarks, the RPO-enhanced version consistently and significantly outperforms its standard counterpart. For instance, R-DPO improves upon DPO by up to +5.4 points on Arena-Hard, and R-IPO surpasses IPO by +7.0 points in win rate on AlpacaEval 2. This universal performance lift validates that RPO is a robust wrapper that can be applied to a wide range of existing preference alignment methods to enhance their resilience to label noise.
@@
While all methods benefit from RPO, the magnitude of the improvement varies, shedding light on the interplay between different loss functions and our robust training scheme. RPO delivers particularly substantial gains to DPO and IPO, especially on the more capable Llama-3 model. This suggests that algorithms with standard loss formulations (like DPO’s cross-entropy) may be more vulnerable to label noise and thus benefit more from RPO’s adaptive re-weighting.
@@
An interesting insight is that the benefits of RPO appear to scale with the capability of the base model. The performance gains are consistently more pronounced on the Llama-3-8B model than on Mistral-7B. For example, the average improvement on the challenging Arena-Hard benchmark is more than double on Llama-3 compared to Mistral. More powerful models are better equipped to capitalize on the cleaner training signal that RPO provides by identifying and down-weighting noisy data.
5.3 Ablation study
We conduct an ablation study to analyze the sensitivity of RPO to two key hyperparameters: the initial annotator reliability, $\eta_{0}$, and the EMA momentum parameter, $\alpha$. All experiments are performed using the R-DPO algorithm on the Mistral-7B-Instruct-v0.2 model. The results are summarized in Table 3.
@@ 5.3 Ablation study
Figure 2: Empirical verification of annotator reliability estimation. Ground-truth reliability ($\eta$ GPT-4o) is established using GPT-4o’s labels. Different reliability levels are simulated by injecting synthetic noise into copies of the UltraFeedback dataset. In the two-annotator setting (b), Annotator 1 is the baseline with no added noise, while noise is progressively added to Annotator 2’s data. The plots confirm that the RPO-estimated reliability ($\eta$ RPO, dashed lines) closely tracks the ground-truth values (solid lines).
@@
The results in Figure 2 show that the estimated reliability $\eta$ (RPO) closely tracks the ground-truth $\eta$ (GPT-4o) in both single-annotator (Figure 2(a)) and two-annotator (Figure 2(b)) settings. Notably, in the two-annotator experiment, RPO successfully identifies the stable reliability of the baseline annotator while accurately tracking the declining reliability of the noisy one.
@@ 6 Conclusion
In this paper, we introduce Robust Preference Optimization (RPO), a novel framework designed to address the critical challenge of aligning LLMs with noisy human preference data. Our approach is distinct from existing methods as it employs an Expectation-Maximization algorithm to infer the reliability of each preference pair, treating labels as soft, dynamic weights rather than fixed ground truths. As a meta-framework, RPO consistently enhances multiple state-of-the-art alignment algorithms, achieving significant performance gains (up to a 7.0% win rate increase on AlpacaEval 2) across various base models.
@@ Appendix A Derivation of general probabilistic model
Here we provide the detailed derivation for (2). For a given prompt $x$ and candidate responses $y_{w},y_{l}$, we assume the probability of the ground-truth preference $y_{w}\succ^{\ast}y_{l}$ is proportional to $\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l}))$. That is:
$p(y_{w}\succ^{\ast}y_{l}|x,\theta)\propto\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l}))$ (8)
@@
$p(y_{l}\succ^{\ast}y_{w}|x,\theta)\propto\exp(-\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w}))$ (9)
@@
$=\sigma\left(\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w})-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l})\right)$
The last line is the General Probabilistic Model in (2).
@@ Appendix B Consistency with Bradley-Terry model for DPO
We show that (2) is consistent with the Bradley-Terry model when applied to DPO. The DPO loss for a preferred pair $(y_{w},y_{l})$ given prompt $x$ is:
$\mathcal{L}_{\text{DPO}}(x,y_{w}\succ y_{l})=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ (10)
@@
Substituting these into our general probabilistic model ((2)):
@@
$=\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
Appendix C Derivation of the RPO EM algorithm
The primary objective of Robust Preference Optimization (RPO) is to find the model parameters $\theta$ and the vector of annotator reliabilities $\boldsymbol{\eta}$ that maximize the log-likelihood of the observed data. The observed data consists of prompts, chosen and rejected responses, and the annotator’s index, denoted as $X=\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i},k_{i})\}_{i=1}^{N}$.
@@ Appendix C Derivation of the RPO EM algorithm
$\mathcal{L}(\theta,\boldsymbol{\eta})=\sum_{i=1}^{N}\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)\eta_{k_{i}}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)(1-\eta_{k_{i}})\right]$ (11)
@@ C.1 Derivation of the Q-Function (The E-Step)
$\mathcal{L}_{c}(\theta,\boldsymbol{\eta};X,Z)=\sum_{i=1}^{N}\left(z_{i}\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)\eta_{k_{i}}\right]+(1-z_{i})\log\left[p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)(1-\eta_{k_{i}})\right]\right)$ (12)
@@
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})\equiv\mathbb{E}_{Z|X,\theta^{(t)},\boldsymbol{\eta}^{(t)}}[\mathcal{L}_{c}(\theta,\boldsymbol{\eta};X,Z)]$ (13)
@@
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})=\sum_{i=1}^{N}\left(\mathbb{E}[z_{i}]\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)\eta_{k_{i}}\right]+(1-\mathbb{E}[z_{i}])\log\left[p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)(1-\eta_{k_{i}})\right]\right)$ (14)
@@
$=\frac{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}}{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})(1-\eta_{k_{i}}^{(t)})}$ (15)
@@
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})=\sum_{i=1}^{N}\left[w_{i}^{(t)}\log(p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)\eta_{k_{i}})+(1-w_{i}^{(t)})\log(p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)(1-\eta_{k_{i}}))\right]$ (16)
C.2 Deriving the RPO Framework (The M-Step)
The goal of the M-Step is to find the parameters for the next iteration, $(\theta^{(t+1)},\boldsymbol{\eta}^{(t+1)})$, by maximizing the Q-function that was constructed using the parameters from the current iteration $t$.
$(\theta^{(t+1)},\boldsymbol{\eta}^{(t+1)})=\arg\max_{\theta,\boldsymbol{\eta}}Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})$ (17)
To perform this maximization, we can first expand the Q-function by separating the terms involving the policy $\theta$ from those involving the annotator reliabilities $\boldsymbol{\eta}$.
@@ C.2 Deriving the RPO Framework (The M-Step)
$+\underbrace{\sum_{i=1}^{N}\left[w_{i}^{(t)}\log\eta_{k_{i}}+(1-w_{i}^{(t)})\log(1-\eta_{k_{i}})\right]}_{\text{Depends only on }\boldsymbol{\eta}}$ (18)
@@
$=\arg\min_{\theta}\left(-\sum_{i=1}^{N}\left[w_{i}^{(t)}\log p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)+(1-w_{i}^{(t)})\log p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)\right]\right)$ (19)
The expression inside the $\arg\min$ is precisely the weighted RPO loss function, $\mathcal{L}_{\text{RPO}}(\theta)$. This establishes that the M-step for the policy parameters is equivalent to minimizing this weighted loss, using weights $w_{i}^{(t)}$ from the E-step.
@@
$\frac{1}{\eta_{k}}\sum_{i:k_{i}=k}w_{i}^{(t)}$ $=\frac{1}{1-\eta_{k}}\left(N_{k}-\sum_{i:k_{i}=k}w_{i}^{(t)}\right)$ (24)

### v2

@@ 5.1 Experimental setup
All experiments were conducted using the Hugging Face TRL library (von Werra et al., 2020). We performed full-parameter fine-tuning for all models, leveraging DeepSpeed ZeRO-3 (Rasley et al., 2020) to manage memory across our hardware setup. Models were trained for a single epoch using the AdamW optimizer (Loshchilov and Hutter, 2019) with a cosine learning rate scheduler, a peak learning rate of $5\times 10^{-7}$, and a global batch size of 64. We also maintained an exponential moving average (EMA) of the model parameters with a decay rate of $\alpha=0.1$. The temperature for DPO and its variants was set to $\beta=0.01$. For all LCPO variants, the reliability parameter was initialized to $\eta^{(0)}=0.9$ and was updated via the E-step after the training epoch was complete. All training was performed on a cluster of 8x NVIDIA A800 40GB GPUs.
@@
To demonstrate that LCPO operates as a versatile meta-framework, we benchmark it against four popular direct preference alignment methods: DPO (Rafailov et al., 2023); IPO (Azar et al., 2023), which uses a squared hinge loss to optimize preferences; SimPO (Meng et al., 2024), which proposes a simplified, reference-free reward formulation normalized by sequence length; and CPO (Xu et al., 2024), which adds a term to directly maximize the likelihood of the preferred response. For each of these baselines, whose loss functions are detailed in Table 1, we compare the original algorithm to its LCPO-enhanced counterpart (e.g., DPO vs. LCPO-DPO).
@@ 5.2 Main results
As shown in Table 2, our experimental results provide strong evidence for the effectiveness of the LCPO framework. We summarize the key findings below.
LCPO as a meta-framework.
The primary takeaway from our experiments is that LCPO functions as a highly effective and generalizable meta-framework. Across all four baseline algorithms (DPO, IPO, SimPO, CPO), both base models (Mistral-7B and Llama-3-8B), and all evaluation benchmarks, the LCPO-enhanced version consistently and significantly outperforms its standard counterpart. For instance, LCPO-DPO improves upon DPO by up to +5.4 points on Arena-Hard, and LCPO-IPO surpasses IPO by +7.0 points in win rate on AlpacaEval 2. This universal performance lift validates that LCPO is a robust wrapper that can be applied to a wide range of existing preference alignment methods to enhance their resilience to label noise.
@@
While all methods benefit from LCPO, the magnitude of the improvement varies, shedding light on the interplay between different loss functions and our robust training scheme. LCPO delivers particularly substantial gains to DPO and IPO, especially on the more capable Llama-3 model. This suggests that algorithms with standard loss formulations (like DPO’s cross-entropy) may be more vulnerable to label noise and thus benefit more from LCPO’s adaptive re-weighting.
@@
An interesting insight is that the benefits of LCPO appear to scale with the capability of the base model. The performance gains are consistently more pronounced on the Llama-3-8B model than on Mistral-7B. For example, the average improvement on the challenging Arena-Hard benchmark is more than double on Llama-3 compared to Mistral. More powerful models are better equipped to capitalize on the cleaner training signal that LCPO provides by identifying and down-weighting noisy data.
@@
Figure 2: Empirical verification of annotator reliability estimation. Ground-truth reliability ($\eta$ GPT-4o) is established using GPT-4o’s labels. Different reliability levels are simulated by injecting synthetic noise into copies of the UltraFeedback dataset. In the two-annotator setting (b), Annotator 1 is the baseline with no added noise, while noise is progressively added to Annotator 2’s data. The plots confirm that the LCPO-estimated reliability ($\eta$ LCPO, dashed lines) closely tracks the ground-truth values (solid lines).
5.3 Ablation study
We conduct an ablation study to analyze the sensitivity of LCPO to two key hyperparameters: the initial annotator reliability, $\eta_{0}$, and the EMA momentum parameter, $\alpha$. All experiments are performed using the LCPO-DPO algorithm on the Mistral-7B-Instruct-v0.2 model. The results are summarized in Table 3.
@@ 5.3 Ablation study
The results in Figure 2 show that the estimated reliability $\eta$ (LCPO) closely tracks the ground-truth $\eta$ (GPT-4o) in both single-annotator (Figure 2(a)) and two-annotator (Figure 2(b)) settings. Notably, in the two-annotator experiment, LCPO successfully identifies the stable reliability of the baseline annotator while accurately tracking the declining reliability of the noisy one.
@@ 6 Conclusion
In this paper, we introduce Latent Collective Preference Optimization (LCPO), a novel framework designed to address the critical challenge of aligning LLMs with noisy human preference data. Our approach is distinct from existing methods as it employs an Expectation-Maximization algorithm to infer the reliability of each preference pair, treating labels as soft, dynamic weights rather than fixed ground truths. As a meta-framework, LCPO consistently enhances multiple state-of-the-art alignment algorithms, achieving significant performance gains (up to a 7.0% win rate increase on AlpacaEval 2) across various base models.
@@ Appendix A Derivation of general probabilistic model
Here we provide the detailed derivation for equation (2). For a given prompt $x$ and candidate responses $y_{w},y_{l}$, we assume the probability of the ground-truth preference $y_{w}\succ^{\ast}y_{l}$ is proportional to $\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l}))$. That is:
$p(y_{w}\succ^{\ast}y_{l}|x,\theta)\propto\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l})).$ (8)
@@
$p(y_{l}\succ^{\ast}y_{w}|x,\theta)\propto\exp(-\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w})).$ (9)
@@
$=\sigma\left(\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w})-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l})\right).$
The last line is the General Probabilistic Model in equation (2).
@@ Appendix B Consistency with Bradley-Terry model for DPO
We show that equation (2) is consistent with the Bradley-Terry model when applied to DPO. The DPO loss for a preferred pair $(y_{w},y_{l})$ given prompt $x$ is:
$\mathcal{L}_{\text{DPO}}(x,y_{w}\succ y_{l})=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right).$ (10)
@@
Substituting these into our general probabilistic model:
@@
$=\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right).$
@@
Appendix C Derivation of the LCPO EM algorithm
The primary objective of Latent Collective Preference Optimization (LCPO) is to find the model parameters $\theta$ and the vector of annotator reliabilities $\boldsymbol{\eta}$ that maximize the log-likelihood of the observed data. The observed data consists of prompts, chosen and rejected responses, and the annotator’s index, denoted as $X=\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i},k_{i})\}_{i=1}^{N}$.
@@ Appendix C Derivation of the LCPO EM algorithm
$\mathcal{L}(\theta,\boldsymbol{\eta})=\sum_{i=1}^{N}\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)\eta_{k_{i}}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)(1-\eta_{k_{i}})\right].$ (11)
@@ C.1 Derivation of the Q-Function (The E-Step)
$\mathcal{L}_{c}(\theta,\boldsymbol{\eta};X,Z)=\sum_{i=1}^{N}\left(z_{i}\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)\eta_{k_{i}}\right]+(1-z_{i})\log\left[p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)(1-\eta_{k_{i}})\right]\right).$ (12)
@@
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})\equiv\mathbb{E}_{Z|X,\theta^{(t)},\boldsymbol{\eta}^{(t)}}[\mathcal{L}_{c}(\theta,\boldsymbol{\eta};X,Z)].$ (13)
@@
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})=\sum_{i=1}^{N}\left(\mathbb{E}[z_{i}]\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)\eta_{k_{i}}\right]+(1-\mathbb{E}[z_{i}])\log\left[p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)(1-\eta_{k_{i}})\right]\right).$ (14)
@@
$=\frac{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}}{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})(1-\eta_{k_{i}}^{(t)})}.$ (15)
@@
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})=\sum_{i=1}^{N}\left[w_{i}^{(t)}\log(p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)\eta_{k_{i}})+(1-w_{i}^{(t)})\log(p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)(1-\eta_{k_{i}}))\right].$ (16)
C.2 Deriving the LCPO Framework (The M-Step)
The goal of the M-Step is to find the parameters for the next iteration, $(\theta^{(t+1)},\boldsymbol{\eta}^{(t+1)})$, by maximizing the Q-function that was constructed using the parameters from the current iteration $t$:
$(\theta^{(t+1)},\boldsymbol{\eta}^{(t+1)})=\arg\max_{\theta,\boldsymbol{\eta}}Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)}).$ (17)
To perform this maximization, we can first expand the Q-function by separating the terms involving the policy $\theta$ from those involving the annotator reliabilities $\boldsymbol{\eta}$:
@@ C.2 Deriving the LCPO Framework (The M-Step)
$+\underbrace{\sum_{i=1}^{N}\left[w_{i}^{(t)}\log\eta_{k_{i}}+(1-w_{i}^{(t)})\log(1-\eta_{k_{i}})\right]}_{\text{Depends only on }\boldsymbol{\eta}}.$ (18)
@@
$=\arg\min_{\theta}\left(-\sum_{i=1}^{N}\left[w_{i}^{(t)}\log p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)+(1-w_{i}^{(t)})\log p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)\right]\right).$ (19)
The expression inside the $\arg\min$ is precisely the weighted LCPO loss function, $\mathcal{L}_{\text{LCPO}}(\theta)$. This establishes that the M-step for the policy parameters is equivalent to minimizing this weighted loss, using weights $w_{i}^{(t)}$ from the E-step.
@@
$\frac{1}{\eta_{k}}\sum_{i:k_{i}=k}w_{i}^{(t)}$ $=\frac{1}{1-\eta_{k}}\left(N_{k}-\sum_{i:k_{i}=k}w_{i}^{(t)}\right),$ (24)